# Quantumness certification via non-demolition measurements

### v1

@@ Abstract
The fundamental question of when a static or dynamic system should be deemed intrinsically quantum remains a challenge to address in absolute terms. A rigorous criterion, however, can be established by focusing on the measurable or reconstructible features of the system. This determination transcends mere issues of a system’s classical simulability or computational complexity. Instead, the critical requirement lies in the certification (ideally, in real-time) of the emergence and persistence of genuine quantum features, principally entanglement and quantum superposition. Quantum Non-Demolition Measurements (QNDM) serve as the appropriate instrument for this certification, both from a theoretical and experimental standpoint. In this review paper, we demonstrate, with accessible clarity, how the implementation of QNDM can be directly linked to a necessary and sufficient condition for the violation of macrorealism in finite-dimensional systems, establishing a conceptual parallel with Leggett-Garg inequalities. Using concrete examples that detail the detection of negative terms in the quasi-probability density function resulting from QNDM, we introduce the core concepts for certifying genuinely quantum features. As specific examples, we discuss an application where the quantum-to-classical transition due to the interaction with an environment can be tracked by QNDM. Moreover, we argue about the robustness of QNDM protocols in the presence of noise sources and their advantages with respect to the Leggett-Garg inequalities. Because of its straightforward implementation, the QNDM approach can be of direct relevance to both the foundations of quantum mechanics and quantum information theory, where a controlled generation and certification of genuinely quantum resources is a central concern.
@@ I Introduction
The main quantum features that cannot be developed by any classical system are entanglement and superposition (or interference). They are related to the concepts of nonlocality and macrorealism (MR), respectively. Traditionally, the criteria used to distinguish between classical and quantum systems and to identify the presence of such features are the famous Bell’s inequalities [1, 2, 3, 4] and the Leggett-Garg inequalities (LGIs) [5]. The former focuses on nonlocality and quantum correlations related to entanglement, while the latter concerns quantum superposition.
@@
In this review, we focus on the macrorealistic properties of quantum systems, starting from the concept of macrorealism introduced by Leggett and Garg in 1985 [5]. According to the latter, a classical system must satisfy three main assumptions: $1)$ Macrorealism per se (MRps), which states that the system is at all times in one of the states available to it; $2)$ Noninvasive Measurability (NIM), which requires that it should be possible to determine the state of a system without disturbing its subsequent dynamics; and $3)$ Induction, added and clarified later [6], which asserts that future measurements cannot influence the present state. The conjunction of these assumptions can be tested experimentally. At a practical level, the NIM condition is difficult to implement. For this reason, it is often replaced by a statistical version called no-signalling in time (NSIT) [5, 7, 8, 9, 10, 6]. More precisely, the NSIT conditions require that the unperturbed marginals of the probability distribution, defined at multiple times, can be recovered, where unperturbed means no changes in the state of the system are induced by quantum measurements.
@@
By measuring time correlations between sequential measurements at two times, Leggett and Garg derived a set of inequalities [5], now known as LGIs, which are formally similar to Bell’s inequalities. If these inequalities are not satisfied, then the system violates the macrorealism (MR) and therefore exhibits quantum features. LGIs, like Bell’s inequalities, provide only a sufficient condition for the violation of MR. Indeed, there are situations in which a system violates MR, i.e., it has quantum properties, while still satisfying the LGIs.
@@
Clearly, a protocol that provides both necessary and sufficient conditions for the violation of MR would be highly advantageous for identifying quantum features. Here, we review some recent works that address this gap by introducing an experimentally feasible and powerful protocol known as Quantum Non-Demolition Measurements (QNDM). This approach builds on measurement schemes proposed at the end of 1970s [11, 12, 13, 14, 15], which have been modified and adapted to extract information from sequential measurements [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], such that the assumptions defining the MR are fulfilled.
@@
The key idea of QNDM is to store the information about a desired observable in the phase of a quantum detector that is coupled to the system. The coupling is designed so that it leaves the statistics of these observables unchanged. Then, after a sequence of system-detector interactions, the phase of the detector is measured, and from it one can reconstruct a quasi-probability density function of the measurement outcomes. The latter is not always positive, and it can be proven [26] that the presence of negative regions is unequivocally associated with the violation of MR. In other words, the quasi-probability density function has negative regions if and only if MRps is violated, and thus the presence of such regions provides both a necessary and sufficient condition for the violation of MR.
Comparable conditions have been presented in the literature [27, 28, 29, 30], but in general they are weaker, namely are satisfied only for specific parameter choices, and are too difficult to verify in practice. In particular, in order to satisfy the conditions in Refs. [27, 28, 29, 30], one has to solve a complex system of parametric equations, whose size grows with a power of the dimension of the measured quantum system. These limitations make the QNDM approach a privileged tool to study and identify the violation of MR and the presence of quantum behaviour.
@@
The article is organized as follows. In Sec. II, we present a physical scenario in which we aim to extract information about work and internal energy variation of a quantum system. This is a prototypical example where the QNDM approach gives a consistent advantage over the standard projective measurements. It also provides the basis for Sec. III, where we introduce the QNDM protocol that is designed to preserve interference effects and quantum superposition. In Sec. IV, we discuss the relation between the QNDM protocol and LGIs, and we prove that it provides a necessary and sufficient condition for the violation of MR. Then, we discuss the resources needed to implement the QNDM compared to the LGIs (Sec.  V) and an example where QNDM can be used to track the quantum-to-classical transition induced an external environment (Sec. VI), showing that the protocol is both robust against noise and efficient in terms of the resources required for its implementation. Finally, Sec. VII contains the conclusions and outlines future perspectives.
@@
II Motivation: Measuring work, heat, and internal energy variation at the quantum level
@@
Work, heat, and internal energy variation are cornerstone concepts in classical thermodynamics and mechanics. It is thus natural to ask how one can define or measure them when the dimension of the system under scrutiny decreases, and quantum effects become relevant. To illustrate the core problem but avoiding irrelevant complexity, we focus on the mechanical definition of internal energy variation and work. In other terms, we consider a system that is isolated from the environment, i.e., it does not dissipate energy in terms of heat, and that is driven by an external force that does work on the system. In this case, the work $\mathcal{W}$ done by the external force is equal to the variation of the internal energy $\Delta\mathcal{U}$: $\mathcal{W}=\Delta\mathcal{U}$.
However, when we reach the quantum regime, we immediately face a problem whenever we need to measure and characterize the statistics of physical quantities intrinsically defined at two or multiple times. Indeed, in classical physics, we implicitly assume that we are able to measure these physical quantities without perturbing the system at any time. Instead, in quantum mechanics, the measurement procedure necessarily perturbs the system by inducing a collapse of its wave function [25].
This problem is well illustrated in the measurement of $\mathcal{W}$ or $\Delta\mathcal{U}$ in a quantum system. From a classical perspective, we can measure sequentially the energy of the system at the beginning $\mathcal{E}_{i}$ and at the end $\mathcal{E}_{f}$ of the driving protocol, and taking their difference: $\mathcal{W}=\Delta\mathcal{U}=\mathcal{E}_{f}-\mathcal{E}_{i}$. Here, all the physical quantities are well defined, and we thus have a clear theoretical definition that matches the measuring procedure.
In the quantum case, we suppose to know the time-dependent Hamiltonian $\hat{H}(t)$ that defines the driving protocol. In particular, the time dependence of $\hat{H}(t)$ is due to the presence of a classical external time-dependent force that exchanges energy with the quantum system, by doing work on it. The system evolves under $\hat{H}(t)$ from time $t=0$ to time $t=\mathcal{T}$, undergoing a unitary transformation, which we denote as $\hat{U}(\mathcal{T})\equiv\hat{U}$. The initial state of the system is supposed to be described by the density operator $\hat{\rho}^{0}$. As in the classical case, if the system is isolated, then the work is equal to the variation of the internal energy, but with a caveat. Indeed, the request is that the initial and final energies of the classical system are substituted by the statistical averages (expectation value) obtained from measuring the quantum system at times $t=0$ and $t=\mathcal{T}$:
@@
Initially, this issue has been circumvented because the first studies considered initializing the quantum system under scrutiny in a thermal state built over the initial Hamiltonian $\hat{H}(0)$, for which sequential measurements suffice to determine two-time correlators. In such cases, one can adopt the so-called Two-Point Measurement (TPM) protocol [36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. This consists of measuring the system energy at the initial time $t=0$, letting the system evolve under the external drive, and then measuring the energy at time $t=\mathcal{T}$. If we denote the eigenvalues and eigenvectors of $\hat{H}(t)$ with $\epsilon_{j}(t)$ and $\mathinner{|{\epsilon_{j}(t)}\rangle}$, respectively, it can be shown [36, 37] that the statistics of the work (or internal energy variation) is given by the probability distribution
@@
where $P_{i}^{(0)}\mathrel{\mathop{\ordinarycolon}}=\langle\epsilon_{i}(0)|\hat{\rho}^{0}|\epsilon_{i}(0)\rangle$ is the probability to find the system initially in the state $\mathinner{|{\epsilon_{i}(0)}\rangle}$, and $P^{(\mathcal{T})}_{i\rightarrow j}\mathrel{\mathop{\ordinarycolon}}=|\langle\epsilon_{j}(\mathcal{T})|\hat{U}|\epsilon_{i}(0)\rangle|^{2}$ is the probability for the occurrence of the transition from $\mathinner{|{\epsilon_{j}(0)}\rangle}$ to $\mathinner{|{\epsilon_{j}(\mathcal{T})}\rangle}$ due to the evolution $\hat{U}$. The average work provided by the TPM protocol can be obtained by the distribution (2) and reads
@@
Eqs. (2) and (3) have an immediate interpretation in simple statistical terms. The work done $\epsilon_{j}(\mathcal{T})-\epsilon_{i}(0)$ during the process is weighted by the probability to start from the state $\mathinner{|{\epsilon_{j}(0)}\rangle}$ and to end in the state $\mathinner{|{\epsilon_{j}(\mathcal{T})}\rangle}$ exactly as it happens in a classical statistical process.
Although its direct application, the first measurement of the TPM protocol causes a perturbation on the system, which leads to the collapse of the initial (at time $t=0$) system’s wave-function. This ultimately entails a change in the dynamics and in the expected prediction [25]. To clarify this point, we consider the definition of the average work as given by Eq. (1) and show that, for a generic initial density operator $\hat{\rho}^{0}$, the TPM protocol cannot recover the full (unperturbed) expression of the average work. Indeed, the right-hand side of Eq. (1) admits two contributions [19]: one related to the diagonal elements of $\hat{\rho}^{0}$, while the other depends on the corresponding off-diagonal elements. Formally, we have:
@@
This observation allows us to make a parallel with the famous two-slit experiment with quantum particles  [46]. If both slits are open, we observe an interference pattern of the particle arriving at the screen. If one of the two slit is closed or we perform a measurement to establish through which slit the particle has passed, the interference pattern vanishes. In the present discussion, the initial projective measurement of $\langle\hat{H}(0)\rangle$ leads to the collapse of the system’s wave-function at $t=0$ and, as a consequence, forces the system into one of the eigenstates of $\langle\hat{H}(0)\rangle$. As in the two-slit experiment, the first measurement of the TPM protocol destroys the interference patterns and changes the subsequent dynamics of the system.
The challenge of measuring the variation of internal energy in a quantum system without destroying the quantum coherence and/or correlations, initially present in the system’s state or generated during its dynamics, has been a subject of recent discussion in the literature. Specifically, pushed by no-go theorems [47, 48] that identify fundamental constraints on measuring physical quantities at two consecutive times, several approaches have been introduced. These include non-linear schemes using information about the initial system’s state [49, 50, 51, 52, 53], initial-time measurements protocols [33, 54], procedures that implement weak measurements [55, 56], and quasi-probability approaches [57, 58, 59, 60, 22, 23, 48, 61, 62, 63, 64, 65]. Among the latter, we count Quantum Non-Demolition Measurements (QNDM) [12, 13, 14, 15], which have been designed in order to acquire information about a quantum system at multiple times while preserving the full quantum dynamics of the system.
@@ III Quantum non-demolition measurement
To better introduce the QNDM scheme, it is preferable to step back for the moment from the specific application discussed in the above section and consider a more general framework [26]. Below, we will show how these concepts directly apply to the measurement of the variation of internal energy.
Let us thus suppose we have a quantum system evolving under a unitary transformation in a time interval $0\leq t\leq\mathcal{T}$. At times $t_{0}\leq t_{1}=\mathcal{T}$, we measure a generic observable $\hat{A}$ (Hermitian operator). In order to realize QNDM, we consider an auxiliary quantum detector to store the desired information in the phase of the detector’s state, which is eventually measured. This scheme allows us to preserve the quantum coherence in the initial density operator of the system, and obtain the quantum correlation function of observables measured at multiple times, without the result being perturbed by the interaction with the detector [20, 21, 31, 22, 23].
Ideally, the sequential non-demolition measurements assume fast system-detector couplings so that, during the interaction with the detector, the system dynamics can be considered frozen. The main point in the QNDM approach is to choose a proper system-detector coupling. If we want to measure the observable $\hat{A}$ we need to implement the unitary transformations $\hat{u}_{k}=\exp\{i(\lambda_{k}/2)\hat{A}\otimes\hat{p}\}$ at any time $t_{k}$ with $k=0,1$ [19, 20, 21]. Here, $\lambda$ is an effective coupling, and $\hat{p}$ is an operator acting on the detector’s degrees of freedom. The unitary transformations $\hat{u}_{k}$ are taken in order to not induce any transition in the system state with respect to the basis that spectrally decomposes $\hat{A}$, whose statistics are thus preserved. Hence, the effect of the system-detector coupling is to make a phase accumulate in the detector state between the eigenstates $\hat{p}$ of $\hat{A}$.
Between the couplings at times $t_{k}$, the system evolves freely from the detector. This means that there is no restriction on the system’s time evolution that can also be driven by an external time-dependent field, which performs work on the system as in Sec. II. Such an evolution from time $t_{0}$ to $t_{1}$ is described by the unitary operator $\hat{U}_{1}\otimes\mathbbm{1}=\hat{U}(t_{1},t_{0})\otimes\mathbbm{1}$ so that the total unitary transformation acting on both the system and detector is $\hat{U}_{tot}=\hat{u}_{1}(\hat{U}_{1}\otimes\mathbbm{1})\hat{u}_{0}$. To keep the notation as simple as possible, we denote $\hat{U}_{i}\equiv\hat{U}_{i}\otimes\mathbbm{1}$ so that $\hat{U}_{tot}$ reads
@@
$\hat{U}_{tot}=\hat{U}_{1}e^{i\frac{\lambda_{1}}{2}\hat{B}\otimes\hat{p}}\,e^{i\frac{\lambda_{0}}{2}\hat{A}\otimes\hat{p}}$. (6)
@@
To understand how the QNDM approach works, it is convenient to decompose the initial state of the system in the basis $\{\mathinner{|{i}\rangle}\}$ in which the observable $\hat{A}$ is diagonal, i.e., $\hat{A}\mathinner{|{i}\rangle}=a_{i}\mathinner{|{i}\rangle}$ with $a_{i}$ real numbers, so that $\mathinner{|{\psi_{0}}\rangle}=\sum_{i}\psi^{0}_{i}\mathinner{|{i}\rangle}$. At the same time, we must initialize the detector state in a coherent superposition of eigenstates of $\hat{p}$, so that it is taken equal to $\mathinner{|{\phi_{0}}\rangle}=\left(1/\sqrt{M}\right)\sum_{p}\mathinner{|{p}\rangle}$. This assures that, during the system-detector coupling, a phase is accumulated in the detector state between the eigenstates $\mathinner{|{p}\rangle}$. Given the initial state $\mathinner{|{\psi_{0}}\rangle}\mathinner{|{\phi_{0}}\rangle}$ and the unitary transformation $\hat{U}_{tot}$, the final system-detector state is (see Ref. [21, 24, 26] for analogous calculation)
$\mathinner{|{\Psi}\rangle}=\frac{1}{\sqrt{M}}\sum_{p}\sum_{i,j}e^{i\frac{\lambda p}{2}(a_{j}+a_{i})}U_{1,ji}\psi^{0}_{i}\mathinner{|{j}\rangle}\mathinner{|{p}\rangle}$ (7)
@@
As the information on the evolution of the monitored system is extracted by measuring the phase accumulated in the detector state (experimentally, it can be done using an interferometric technique [66, 67, 62]), it is worth determining a formal expression for such a phase, whose meaning we are also going to discuss. Denoting with $\hat{R}={\mathinner{|{\Psi}\rangle\langle{\Psi}|}}$, the total density matrix after the evolution, we first trace out the system degrees of freedom to obtain the final density matrix $\hat{r}={\rm Tr}_{S}\left[\hat{R}\right$] of the detector. Second, we take the accumulated phase

### v2

@@ Abstract
The fundamental question of when a static or dynamic system should be deemed intrinsically quantum remains a challenge to address in absolute terms. In this regard, a critical requirement lies in the certification (ideally, in real-time) of the emergence and persistence of genuine quantum features, principally entanglement and quantum superposition. Quantum Non-Demolition Measurements (QNDM) serve as the appropriate instrument for this certification, both from a theoretical and experimental standpoint. In this review paper, we explain, with accessible clarity, how the implementation of QNDM can be directly linked to a necessary and sufficient condition for the presence of genuinely quantum features in the system’s state monitored over time in finite-dimensional systems, establishing a conceptual parallel with Leggett-Garg inequalities. Using concrete examples that detail the detection of negative terms in the quasi-probability density function resulting from QNDM, we introduce the core concepts for quantumness certification. As specific examples, we discuss an application where the quantum-to-classical transition due to the interaction with an environment can be tracked by QNDM. Moreover, we argue about the robustness of QNDM protocols in the presence of noise sources and their advantages with respect to standard Leggett-Garg inequalities defined by two-time correlators.
@@ I Introduction
The main quantum features that cannot be developed by any classical system are entanglement and superposition (or interference). They are related to the concepts of nonlocality and macrorealism (MR), respectively. Traditionally, the criteria used to distinguish between classical and quantum systems, thus identifying the presence of entanglement and quantum superposition, are the famous Bell’s inequalities [1, 2, 3, 4] and the Leggett-Garg inequalities (LGIs) [5]. The former focuses on nonlocality and quantum correlations related to entanglement, while the latter concerns quantum superposition.
@@
Macrorealism is a concept introduced by Leggett and Garg in 1985 [5], whereby a classical system must satisfy three main assumptions: $1)$ Macrorealism per se (MRps), which states that the system is at all times in one of the states available to it; $2)$ Noninvasive Measurability (NIM), which requires that it should be possible to determine the state of a system without disturbing its state and thus its subsequent dynamics; and $3)$ Induction, added and clarified later [6], which asserts that future measurements cannot influence the present state.
@@
By measuring time correlations between sequential measurements at two times, Leggett and Garg derived the set of inequalities [5] known as LGIs, which are formally similar to Bell’s inequalities. If these inequalities are not satisfied, then the system violates the MR and therefore exhibits quantum features. LGIs, like Bell’s inequalities, provide only a sufficient condition for the violation of MR. Indeed, there are situations in which a system violates MR, i.e., it has quantum properties, while still satisfying the LGIs. The conjunction of the MR’s assumptions 1)-3) listed above can be tested experimentally, in strict relation to observation of LGIs’ violations (see, for example Refs. [7, 8, 9, 10]).
@@
In this review, we examine some recent studies that have been developed to address this issue by introducing an experimentally feasible and powerful protocol known as Quantum Non-Demolition Measurements (QNDM). This approach builds on measurement schemes proposed at the end of 1970s [11, 12, 13, 14, 15], which have been modified and adapted to extract information from sequential non-projective measurements [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26].
@@
The key idea of QNDM is to store the information about a desired observable in the phase of a quantum detector that is coupled to the system. The coupling is designed so that it leaves the statistics of these observables unchanged. Then, after a sequence of system-detector interactions, the phase of the detector is measured, and from it one can reconstruct a quasi-probability density function of the measurement outcomes. The latter is not always positive, and it can be proven [26] that the presence of negative regions is unequivocally associated with the presence of genuinely quantum features in the system’s state. More precisely, the quasi-probability density function has negative regions if and only if MRps is violated, namely, at least a quantum superposition of the observable’s eigenstates is identified in the system’s state at the time instants when system-detector interaction occurs. On the other hand, if we find that the quasi-probability density function resulting from QNDM is positive, then we can show that such a distribution can be obtained from sequential projective measurements, and thus described by the so-called Two-Point Measurement (TPM) protocol [27, 28, 29, 30]. It is known that the results provided by the TPM scheme have a classical-quantum correspondence [31, 25, 26].
Similar ideas of using an additional quantum system to reconstruct features of a quantum system have also been proposed in the field of full-counting statistics [32, 33, 34], where the statistics of electrons flowing through a wire is measured. These results led to several implementations in condensed matter and electronic circuits [35, 36, 37], bringing advancement in such fields. While it was known that the approach of full-counting statistics entails a quasi-probability distribution [34, 38, 39, 40, 41], implications in terms of macrorealism violation and the connection of the latter to the negativity of the quasi-probability distribution have not been fully understood.
@@
The article is organized as follows. In Sec. II, we present a physical scenario in which we aim to extract information about work and internal energy variation of a quantum system. This is a prototypical example where the QNDM approach gives a consistent advantage over the standard projective measurements. It also provides the basis for Sec. III, where we introduce the QNDM protocol that is designed to preserve interference effects and quantum superposition. In Sec. IV, we discuss the main properties of the QNDM protocol and its relation with the LGIs, and we quantify the resources needed for their implementation (Sec. V). Then, we report an example where QNDM can be used to track the quantum-to-classical transition induced by an external environment (Sec. VI), showing that the protocol is both robust against noise and efficient in terms of resources. Finally, Sec. VII contains the conclusions and outlines future perspectives.
@@
II Motivation: Describing internal energy variation and work at the quantum level
@@
Work, heat, and internal energy variation are cornerstone concepts in classical thermodynamics and mechanics. It is thus natural to ask how one can define or measure them when the dimension of the system under scrutiny decreases, and quantum effects become relevant. Let us focus on the mechanical definition of internal energy variation and work by considering a system that is isolated from the environment (i.e., it does not dissipate energy in terms of heat) and is driven by an external force that does work on the system. In such a case, the work $\mathcal{W}$ done by the external force is equal to the variation of the internal energy $\Delta\mathcal{U}$: $\mathcal{W}=\Delta\mathcal{U}$.
However, when we reach the quantum regime, we face a problem whenever we need to measure and characterize the statistics of physical quantities defined at two or multiple times. In classical physics, we implicitly assume that we are able to measure these physical quantities without perturbing the system at any time. Instead, in quantum mechanics, the measurement procedure necessarily perturbs the system by inducing a collapse of its wave function [25].
From a classical perspective, we can measure sequentially the energy of the system at the beginning $\mathcal{E}_{i}$ and at the end $\mathcal{E}_{f}$ of the driving protocol, and taking their difference: $\mathcal{W}=\Delta\mathcal{U}=\mathcal{E}_{f}-\mathcal{E}_{i}$. All the physical quantities are well defined, and we thus have a clear theoretical definition that matches the measuring procedure.
In the quantum case, we suppose to know the time-dependent Hamiltonian $\hat{H}(t)$ that defines the driving protocol. The time dependence of $\hat{H}(t)$ is due to the presence of a classical external time-dependent force that exchanges energy with the quantum system, by doing work on it. The system evolves under $\hat{H}(t)$ from time $t=0$ to time $t=\mathcal{T}$, undergoing a unitary transformation, which we denote as $\hat{U}(\mathcal{T})\equiv\hat{U}$. The initial state of the system is supposed to be described by the density operator $\hat{\rho}^{0}$. As in the classical case, if the system is isolated, then the work is equal to the variation of the internal energy, but with a caveat. Indeed, the request is that the initial and final energies of the classical system are substituted by the statistical averages (expectation value) obtained from measuring the quantum system at times $t=0$ and $t=\mathcal{T}$:
@@
Initially, this issue has been circumvented because the first studies on a driven quantum system considered initializing the system in a thermal state built over the initial Hamiltonian $\hat{H}(0)$, for which sequential measurements suffice to determine two-time correlators. In such cases, one can adopt the TPM protocol [46, 47, 27, 28, 29, 30, 48, 49, 50, 51]. This consists of measuring the system energy at the initial time $t=0$, letting the system evolve under the external drive, and then measuring the energy at time $t=\mathcal{T}$. If we denote the eigenvalues and eigenvectors of $\hat{H}(t)$ with $\epsilon_{j}(t)$ and $\mathinner{|{\epsilon_{j}(t)}\rangle}$, respectively, it can be shown [46, 47] that the statistics of the work (or internal energy variation) is given by the probability distribution
@@
where $P_{i}^{(0)}\mathrel{\mathop{\ordinarycolon}}=\langle\epsilon_{i}(0)|\hat{\rho}^{0}|\epsilon_{i}(0)\rangle$ is the probability to find the system initially in the state $\mathinner{|{\epsilon_{i}(0)}\rangle}$, and $P^{(\mathcal{T})}_{i\rightarrow j}\mathrel{\mathop{\ordinarycolon}}=|\langle\epsilon_{j}(\mathcal{T})|\hat{U}|\epsilon_{i}(0)\rangle|^{2}$ is the probability for the occurrence of the transition from $\mathinner{|{\epsilon_{j}(0)}\rangle}$ to $\mathinner{|{\epsilon_{j}(\mathcal{T})}\rangle}$ due to the unitary transformation $\hat{U}$. The average work provided by the TPM protocol can be obtained by the distribution (2) and reads as
@@
Eqs. (2) and (3) have an immediate interpretation in simple statistical terms. The work done $\epsilon_{j}(\mathcal{T})-\epsilon_{i}(0)$ during the process is weighted by the probability to start from the state $\mathinner{|{\epsilon_{j}(0)}\rangle}$ and to end in the state $\mathinner{|{\epsilon_{j}(\mathcal{T})}\rangle}$, exactly as it happens in a classical statistical process.
Although its direct application, the first measurement of the TPM protocol causes a perturbation on the system, which leads to the collapse of the initial (at time $t=0$) system’s wave-function. This ultimately entails a change in the dynamics and in the expected prediction [25]. To clarify this point, we consider the definition of the average work as given by Eq. (1) and show that, for a generic non-diagonal initial density operator $\hat{\rho}^{0}$, the TPM protocol cannot recover the full (unperturbed) expression of the average work. Indeed, the right-hand side of Eq. (1) admits two contributions [19]: one related to the diagonal elements of $\hat{\rho}^{0}$, while the other depends on the corresponding off-diagonal elements. Formally, we have:
@@
This observation allows us to make a parallel with the famous two-slit experiment with quantum particles  [52]. If both slits are open, we observe an interference pattern of the particle arriving at the screen. If one of the two slits is closed or we perform a measurement to establish through which slit the particle has passed, the interference pattern vanishes. In the present discussion, the initial projective measurement of $\langle\hat{H}(0)\rangle$ leads to the collapse of the system’s wave-function at $t=0$ and, as a consequence, forces the system into one of the eigenstates of $\langle\hat{H}(0)\rangle$. As in the two-slit experiment, the first measurement of the TPM protocol destroys the interference patterns and changes the subsequent dynamics of the system.
The challenge of measuring the variation of internal energy in a quantum system without destroying the quantum coherence and/or correlations, initially present in the system’s state or generated during its dynamics, has been a subject of recent discussions in the literature. Specifically, pushed by no-go theorems [53, 54] that identify fundamental constraints on measuring physical quantities at two consecutive times, several approaches have been introduced. These include non-linear schemes using information about the initial system’s state [55, 56, 57, 58, 59], inference protocols with initial-time measurement [43, 60], procedures implementing weak measurements [39, 61], and quasi-probability approaches [62, 63, 64, 65, 22, 23, 54, 66, 67, 68, 69, 70]. Among the latter, we count Quantum Non-Demolition Measurements (QNDM) [12, 13, 14, 15] that have been specifically designed to acquire information about a quantum system at multiple times while preserving the full quantum dynamics of the system.
@@ III Quantum non-demolition measurement
To introduce the QNDM scheme, it is preferable to step back from a specific application and consider a more general framework [26]. Subsequently, we will show how this framework applies to the measurement of the internal energy variation.
Let us thus suppose we have a quantum system evolving under a unitary transformation in a time interval $t_{0}\leq t\leq\mathcal{T}$. At times $t_{0}\leq t_{1}=\mathcal{T}$, we measure a generic observable $\hat{A}$ (Hermitian operator). In order to realize QNDM, we consider an auxiliary quantum detector to store the desired information in the phase of the detector’s state, which is eventually measured. This scheme allows us to preserve the quantum coherence in the initial density operator of the system, and derive the quantum correlation function of observables measured at multiple times, without the result being perturbed by the interaction with the detector [20, 21, 42, 22, 23].
Ideally, the sequential non-demolition measurements assume fast system-detector couplings so that, during the interaction with the detector, the system dynamics can be considered frozen. The main point in the QNDM approach is to choose a proper system-detector coupling. If we want to measure the observable $\hat{A}$ we need to implement the unitary transformations $\hat{u}_{k}=\exp\{i(\lambda_{k}/2)\hat{A}\otimes\hat{p}\}$ at any time $t_{k}$ with $k=0,1$ [19, 20, 21]. Here, $\lambda$ is an effective coupling, and $\hat{p}$ is an operator acting on the detector’s degrees of freedom. The unitary transformations $\hat{u}_{k}$ are taken in order to not induce any transition in the system state with respect to the basis that spectrally decomposes $\hat{A}$, whose statistics are thus preserved. Hence, the effect of the system-detector coupling is to make a phase accumulate in the detector’s state between the eigenstates $\hat{p}$ of $\hat{A}$.
Between the couplings at times $t_{k}$, the system evolves freely from the detector. This means that there is no restriction on the system’s time evolution that can also be driven by an external time-dependent field, which performs work on the system as in Sec. II. Such an evolution from time $t_{0}$ to $t_{1}$ is described by the unitary operator $\hat{U}_{1}\otimes\mathbbm{1}=\hat{U}(t_{1},t_{0})\otimes\mathbbm{1}$ so that the total unitary transformation acting on both the system and detector is $\hat{U}_{tot}=\hat{u}_{1}(\hat{U}_{1}\otimes\mathbbm{1})\hat{u}_{0}$. To keep the notation as simple as possible, we denote $\hat{U}_{i}\equiv\hat{U}_{i}\otimes\mathbbm{1}$ so that $\hat{U}_{tot}$ reads as
@@
$\hat{U}_{tot}=\hat{U}_{1}e^{i\frac{\lambda_{1}}{2}\hat{B}\otimes\hat{p}}\,e^{i\frac{\lambda_{0}}{2}\hat{A}\otimes\hat{p}}$, (6)
@@
To understand how the QNDM approach works, it is convenient to decompose the initial state of the system in the basis $\{\mathinner{|{i}\rangle}\}$ in which the observable $\hat{A}$ is diagonal, i.e., $\hat{A}\mathinner{|{i}\rangle}=a_{i}\mathinner{|{i}\rangle}$ with $a_{i}$ real numbers, so that $\mathinner{|{\psi_{0}}\rangle}=\sum_{i}\psi^{0}_{i}\mathinner{|{i}\rangle}$. At the same time, we must initialize the detector’s state in a coherent superposition of eigenstates of $\hat{p}$, so that it is taken equal to $\mathinner{|{\phi_{0}}\rangle}=\left(1/\sqrt{M}\right)\sum_{p}\mathinner{|{p}\rangle}$. This assures that, during the system-detector coupling, a phase is accumulated in the state of the detector between the eigenstates $\mathinner{|{p}\rangle}$. Given the initial state $\mathinner{|{\psi_{0}}\rangle}\mathinner{|{\phi_{0}}\rangle}$ and the unitary transformation $\hat{U}_{tot}$, the final system-detector’s state is (see Ref. [21, 24, 26] for analogous calculation)
$\mathinner{|{\Psi}\rangle}=\frac{1}{\sqrt{M}}\sum_{p}\sum_{i,j}e^{i\frac{\lambda p}{2}(a_{j}+a_{i})}U_{1,ji}\psi^{0}_{i}\mathinner{|{j}\rangle}\mathinner{|{p}\rangle},$ (7)
@@
As the information on the evolution of the monitored system is extracted by measuring the phase accumulated in the detector’s state (experimentally, it can be done using an interferometric technique [71, 72, 67]), it is worth determining a formal expression for such a phase, whose meaning we are going to discuss. Denoting with $\hat{R}={\mathinner{|{\Psi}\rangle\langle{\Psi}|}}$ the total density matrix after the evolution, we trace out the system’s degrees of freedom to obtain the final density matrix $\hat{r}={\rm Tr}_{S}\left[\hat{R}\right$] of the detector. Then, we take the accumulated phase